(D^4 +13D^2 + 36)y = 0

1
Expert's answer
2022-02-17T04:31:36-0500

Characteristic (auxiliary) equation


"r^4+13r^2+36=0"

"r^2=-4\ or\ r^2=-9"

"r_1=2i, r_2=-2i, r_3=3i, r_4=-3i"

The general solution of the given differential equation is


"y=c_1\cos(2x)+c_2\sin(2x)+c_3\cos (3x)+c_4\sin (3x)"